a drier is used to dry 150kg of salmon for the production of salted fish. after cleaning and cutting, 9% of this went out as waste. if the initial moisture content of fish is 73% and the moisture after drying is 42%, calculate the weight of salted salmon that will be obtained
Answer #1

Total weight of fish is 150 kg. 9% of this went out as waste therefore weight after cleaning and cutting = 150 - 150 x 9/100 = 136.5 kg.

Initial moisture content is 73% therefore total weight of moisture = 136.5 x 73/100 = 99.645 kg and thus total weight of moisture free (complete dry) fish = 136.5 - 99.6 = 36.9 kg.

After drying, the moisture content is 42% therefore total weight of salted salmon after drying is 36.9 + 36.9 x 42/100 = 52.4 kg.